# The Muckenhoupt condition

### v2

@@ 5.3. Musielak-Orlicz spaces
We call a function $\phi:\mathbf{R}^{d}\times[0,\infty)\to[0,\infty]$ a Young function if
@@
$\phi^{\ast}(x,t):=\sup_{u\geq 0}ut-\phi(x,u).$
Given two Young functions $\phi,\psi$, we write $\phi\lesssim\psi$ if there are $\lambda,K>0$ and a $0\leq h\in L^{1}(\mathbf{R}^{d})$ such that for a.e. $x\in\mathbf{R}^{d}$ and all $t\geq 0$
@@
is a Young function satisfying
@@
We say that a Young function $\phi$ satisfies the $\Delta_{2}$ condition (as was originally introduced in [Mus83]), if there exists a constant $K>0$ and a function $0\leq h\in L^{1}(\mathbf{R}^{d})$ such that for a.e. $x\in\mathbf{R}^{d}$ and all $t\geq 0$ we have
@@
Often, the $\Delta_{2}$ condition is defined through this inequality only when $h=0$. By passing to an equivalent Young function, we may assume without loss of generality that this is the case. Indeed, assuming $K>1$, we can set
@@
as desired. If both $\phi$ and $\phi^{\ast}$ satisfy the $\Delta_{2}$ condition, we can also find a Young function $\psi\sim\phi$ for which $\psi$ and $\psi^{\ast}$ satisfy the $\Delta_{2}$ condition both with $h=0$, see [HH19, Theorem 2.5.24].
@@
Let $\phi$ be a Young function. Then the following are equivalent:
@@
Let $\phi$ be a young function for which both $\phi$ and $\phi^{\ast}$ satisfy the $\Delta_{2}$ condition. Then $M:L^{\phi(\cdot)}(\mathbf{R}^{d})\to L^{\phi(\cdot)}(\mathbf{R}^{d})$ if and only if $M:L^{\phi^{\ast}(\cdot)}(\mathbf{R}^{d})\to L^{\phi^{\ast}(\cdot)}(\mathbf{R}^{d})$.
@@
Given a Young function $\phi$, we set
@@
for a.e. $x\in\mathbf{R}^{d}$. We say that $\phi$ satisfies the $(A1)$ condition if there is a $10<\beta<$ 1 for which for every ball $B\subseteq\mathbf{R}^{d}$ with $|B|\leq 1$ we have
@@
To understand the conditions $(A0)$$(A2)$, we remark on which conditions they correspond to in the variable Lebesgue case.
@@
Let $\phi$ be a Young function for which $\phi^{\ast}$ satisfies the $\Delta_{2}$ condition, and $\phi$ satisfies the conditions $(A0)$$(A2)$. Then
@@
Let $\phi$ be a Young function for which $\phi^{\ast}$ satisfies the $\Delta_{2}$ condition, and $\phi$ satisfies $(A0)$$(A2)$. If $w$ is a weight for which $L^{\phi_{w}(\cdot)}(\mathbf{R}^{d})\in A$, then
@@
Let $\phi$ be a young function satisfying conditions $(A0)$$(A2)$. Is it true that $L^{\phi(\cdot)}(\mathbf{R}^{d})\in\mathcal{G}$?
@@
To end this subsection, we comment on the results without the assumptions $(A0)$-$(A2)$ by Diening in [Die05]. Just like in the variable Lebesgue case, he considered the condition $L^{\phi(\cdot)}(\mathbf{R}^{d})\in A_{\text{strong}}$ for $N$-functions, i.e., Young function $\phi$ satisfying the additional properties that for a.e. $x\in\mathbf{R}^{d}$ we have
@@
These conditions exclude, e.g., the function $\phi(x,t):=\phi_{p}(t)$ for $p=1$ or $p=\infty$. While he was unable to prove a characterization of the boundedness of $M$ in $L^{\phi(\cdot)}(\mathbf{R}^{d})$, partial results can be found in [Die05, Theorem 5.7, Theorem 6.4]. In particular, he establishes that $L^{\phi(\cdot)}(\mathbf{R}^{d})\in A_{\text{strong}}$ implies a certain reverse Hölder estimate that generalizes the reverse Hölder estimate for Muckenhoupt weights $w\in A_{p}$. This leaves us with the question of whether or not $L^{\phi(\cdot)}(\mathbf{R}^{d})\in A_{\text{strong}}$ implies the existence of an $r>1$ for which also $L^{\phi(\cdot)}(\mathbf{R}^{d})^{r}\in A_{\text{strong}}$, see Question G. We note here that
@@
where $\phi_{r}(x,t):=\phi(x,t^{\frac{1}{r}})$.
@@ 5.4. Morrey spaces
Let $1<p<q<\infty$. It is shown in [Tan15, Proposition 4.2] that $M$ is bounded on $M_{w}^{p,q}(\mathbf{R}^{d})$ with power weights $w(x)=|x|^{\alpha d}$ if and only if $-\tfrac{1}{q}\leq\alpha<\tfrac{1}{q^{\prime}}$. In particular, this holds in the case $\alpha=-\tfrac{1}{q}$, i.e., for
@@
As noted in the above example, Tanaka showed in [Tan15] that for any power weight $w(x)=|x|^{\alpha d}$ with with $-\tfrac{1}{q}\leq\alpha<\tfrac{1}{q^{\prime}}$, we have $M:M^{p,q}_{w}(\mathbf{R}^{d})\to M^{p,q}_{w}(\mathbf{R}^{d})$. Moreover, he showed that in this case, this is precisely the Muckenhoupt condition $M^{p,q}_{w}(\mathbf{R}^{d})\in A$.
@@
when $w(x):=|x|^{-\frac{d}{q}}$, which leaves open the question of whether this bound is true for the power weights in the $A_{q}$ range, i.e., for $-\tfrac{1}{q}<\alpha<\tfrac{1}{q^{\prime}}$. It was shown in [Nie23, Theorem 3.7] that (5.1) holds for any weight $w$ satisfying
@@
$-\frac{1}{q}<\alpha<\frac{1}{p^{\prime}}.$
@@
Let $1<p\leq q<\infty$ and let $w(x):=|x|^{\alpha d}$ for $-\tfrac{1}{q}<\alpha<\frac{1}{q^{\prime}}$. Then we have the following assertions:
@@
$[w]_{p,\text{loc}}:=\sup_{B}\langle w\rangle_{p,B}\langle w^{-1}\rangle_{p^{\prime},B},$
@@
When $\{x_{n}:n\geq 1\}=\{0\}$, the collection $\mathcal{P}=:\mathcal{P}_{0}$ is the collection of all cubes centered at 00, and this result was already proven by Duoandikoetxea and Rosenthal in [DR21, Theorem 6.1]. As a matter of fact, in [DR22, Theorem 4.1] they showed that all Calderón-Zygmund operators are bounded on $M^{p,q}_{w}(\mathbf{R}^{d};\mathcal{P}_{0})$ if and only if
@@
Combining this with Theorem 1.2, we obtain the following characterization.
@@
Let $\mathcal{P}_{0}$ denote the collection of cubes centered at 00 and let $1<p\leq q<\infty$. Then the following are equivalent:
@@ Appendix A Weak-type bounds for sparse operators
Let $\mathcal{D}$ be a dyadic grid, let $10<\eta<$ 1, and let $\mathcal{S}\subseteq\mathcal{D}$ be a finite $\eta$-sparse collection. Then for all $10<\nu<$ 1 and all $f\in L^{1}_{\text{loc}}(\mathbf{R}^{d})$ there exists a collection $\mathcal{E}\subseteq\mathcal{D}$ such that
@@
In particular, $\mathcal{E}$ is $\nu$-sparse with $E_{Q}:=\bigcup_{Q^{\prime}\in\text{ch}_{\mathcal{E}}(Q)}Q^{\prime}$.
@@
and $\widehat{Q}\backslash E\neq\varnothing$ for all $Q\in\text{ch}_{\mathcal{E}}(Q_{0})$, where $\widehat{Q}$ denotes the dyadic parent of $Q$. Then
@@
Fix $Q\in\text{ch}_{\mathcal{E}}(Q_{0})$ and pick $\widehat{x}\in\widehat{Q}\backslash E$. Then
@@
Let $\mathcal{D}$ be a dyadic grid, let $10<\nu<$ 1, and let $\mathcal{S}\subseteq\mathcal{D}$ be a finite collection satisfying
@@
$\int_{\{A_{\mathcal{S}}f>2\}\backslash\{M^{\mathcal{D}}f>\frac{1}{4}\}}\!|g|\,\mathrm{d}x\leq\sum_{m=1}^{\infty}4^{-m}\sum_{Q\in\mathcal{S}_{m}}\int_{F_{m}(Q)}\!|g|\,\mathrm{d}x.$
@@
Now fix $m\geq 1$. We let $\mathcal{S}_{m,0}$ denote the maximal cubes (with respect to inclusion) in $\mathcal{S}_{m}$. Note that these exist, as $\mathcal{S}$ was assumed to be finite. Moreover, we iteratively define $\mathcal{S}_{m,n}$ as the maximal cubes in $\mathcal{S}_{m}\backslash\bigcup_{k=0}^{n-1}\mathcal{S}_{m,k}$. Again since $\mathcal{S}$ is finite, this means that there is an $N\geq 0$ for which
@@
For $Q\in\mathcal{S}_{m,n}$ we set $E(Q):=Q\backslash\bigcup_{Q^{\prime}\in\mathcal{S}_{m,n+1}}Q^{\prime}$ so that the collection $(E(Q))_{Q\in\mathcal{S}_{m}}$ is pairwise disjoint.

### v3

@@ 5.3. Musielak-Orlicz spaces
We call a function $\phi:\mathbf{R}^{d}\times[0,\infty)\to[0,\infty]$ a generalized $\Phi$-function if
@@
$\phi^{\ast}(x,t):=\sup_{u\geq 0}\{ut-\phi(x,u)\}.$
Given two generalized $\Phi$-functions $\phi,\psi$, we write $\phi\lesssim\psi$ if there are $\lambda,K>0$ and a $0\leq h\in L^{1}(\mathbf{R}^{d})$ such that for a.e. $x\in\mathbf{R}^{d}$ and all $t\geq 0$
@@
is a generalized $\Phi$-function satisfying
@@
We say that a generalized $\Phi$-function $\phi$ satisfies the $\Delta_{2}$ condition (as was originally introduced in [Mus83]), if there exists a constant $K>0$ and a function $0\leq h\in L^{1}(\mathbf{R}^{d})$ such that for a.e. $x\in\mathbf{R}^{d}$ and all $t\geq 0$ we have
@@
Often, the $\Delta_{2}$ condition is defined through this inequality only when $h=0$. By passing to an equivalent generalized $\Phi$-function, we may assume without loss of generality that this is the case. Indeed, assuming $K>1$, we can set
@@
as desired. If both $\phi$ and $\phi^{\ast}$ satisfy the $\Delta_{2}$ condition, we can also find a generalized $\Phi$-function $\psi\sim\phi$ for which $\psi$ and $\psi^{\ast}$ satisfy the $\Delta_{2}$ condition both with $h=0$, see [HH19, Theorem 2.5.24].
@@
Let $\phi$ be a generalized $\Phi$-function. Then the following are equivalent:
@@
Let $\phi$ be a generalized $\Phi$-function for which both $\phi$ and $\phi^{\ast}$ satisfy the $\Delta_{2}$ condition. Then $M:L^{\phi(\cdot)}(\mathbf{R}^{d})\to L^{\phi(\cdot)}(\mathbf{R}^{d})$ if and only if $M:L^{\phi^{\ast}(\cdot)}(\mathbf{R}^{d})\to L^{\phi^{\ast}(\cdot)}(\mathbf{R}^{d})$.
@@
Given a generalized $\Phi$-function $\phi$, we set
@@
for a.e. $x\in\mathbf{R}^{d}$. We say that $\phi$ satisfies the $(A1)$ condition if there is a $0<\beta<1$ for which for every ball $B\subseteq\mathbf{R}^{d}$ with $|B|\leq 1$ we have
@@
To understand the conditions (A0)(A0)$(A2)$, we remark on which conditions they correspond to in the variable Lebesgue case.
@@
Let $\phi$ be a generalized $\Phi$-function for which $\phi^{\ast}$ satisfies the $\Delta_{2}$ condition, and $\phi$ satisfies the conditions (A0)(A0)$(A2)$. Then
@@
Let $\phi$ be a generalized $\Phi$-function for which $\phi^{\ast}$ satisfies the $\Delta_{2}$ condition, and $\phi$ satisfies (A0)(A0)$(A2)$. If $w$ is a weight for which $L^{\phi_{w}(\cdot)}(\mathbf{R}^{d})\in A$, then
@@
Let $\phi$ be a generalized $\Phi$-function satisfying conditions (A0)(A0)$(A2)$. Is it true that $L^{\phi(\cdot)}(\mathbf{R}^{d})\in\mathcal{G}$?
@@
To end this subsection, we comment on the results without the assumptions (A0)(A0)-$(A2)$ by Diening in [Die05]. Just like in the variable Lebesgue case, he considered the condition $L^{\phi(\cdot)}(\mathbf{R}^{d})\in A_{\text{strong}}$ for an $N$-function $\phi$, i.e., a generalized $\Phi$-function satisfying the additional properties that for a.e. $x\in\mathbf{R}^{d}$ we have
@@
These conditions exclude, e.g., the function $\phi(x,t):=\phi_{p}(t)$ for $p=1$ or $p=\infty$. While he was unable to prove a characterization of the boundedness of $M$ in $L^{\phi(\cdot)}(\mathbf{R}^{d})$, he does present partial results.
In [Die05, Theorem 5.7] he establishes that $L^{\phi(\cdot)}(\mathbf{R}^{d})\in A_{\text{strong}}$ is equivalent to a certain reverse Hölder estimate that generalizes the reverse Hölder estimate for Muckenhoupt weights $w\in A_{p}$. This leaves us with the question of whether or not $L^{\phi(\cdot)}(\mathbf{R}^{d})\in A_{\text{strong}}$ implies the existence of an $r>1$ for which also $L^{\phi(\cdot)}(\mathbf{R}^{d})^{r}\in A_{\text{strong}}$, see Question G. We note here that
@@
where $\phi_{r}(x,t):=\phi(x,t^{\frac{1}{r}})$. Moreover, in [Die05, Theorem 6.4] he shows that if a certain stronger reverse Hölder condition holds, then there is an $r>1$ for which $M$ is bounded on $L^{\phi(\cdot)}(\mathbf{R}^{d})^{r}$.
@@ 5.4. Morrey spaces
Let $1<p<q<\infty$. It is shown in [Tan15, Proposition 4.2] that $M$ is bounded on $M_{w}^{p,q}(\mathbf{R}^{d})$ with power weights $w(x)=|x|^{\alpha d}$ if and only if $\tfrac{1}{q}\leq\alpha<\tfrac{1}{q^{\prime}}$. In particular, this holds in the case $\alpha=-\tfrac{1}{q}$, i.e., for
@@
As noted in the above example, Tanaka showed in [Tan15] that for any power weight $w(x)=|x|^{\alpha d}$ with with $\tfrac{1}{q}\leq\alpha<\tfrac{1}{q^{\prime}}$, we have $M:M^{p,q}_{w}(\mathbf{R}^{d})\to M^{p,q}_{w}(\mathbf{R}^{d})$. Moreover, he showed that in this case, this is precisely the Muckenhoupt condition $M^{p,q}_{w}(\mathbf{R}^{d})\in A$.
@@
when $w(x):=|x|^{-\frac{d}{q}}$, which leaves open the question of whether this bound is true for the power weights in the $A_{q}$ range, i.e., for $\tfrac{1}{q}<\alpha<\tfrac{1}{q^{\prime}}$. It was shown in [Nie23, Theorem 3.7] that (5.1) holds for any weight $w$ satisfying
@@
$\frac{1}{q}<\alpha<\frac{1}{p^{\prime}}.$
@@
Let $1<p\leq q<\infty$ and let $w(x):=|x|^{\alpha d}$ for $\tfrac{1}{q}<\alpha<\frac{1}{q^{\prime}}$. Then we have the following assertions:
@@
$w]_{p,\text{loc}}:=\sup_{B}\langle w\rangle_{p,B}\langle w^{-1}\rangle_{p^{\prime},B},$
@@
When $\{x_{n}:n\geq 1\}=\{0\}$, the collection $\mathcal{P}=:\mathcal{P}_{0}$ is the collection of all cubes centered at 0, and this result was already proved by Duoandikoetxea and Rosenthal in [DR21, Theorem 6.1]. As a matter of fact, in [DR22, Theorem 4.1] they showed that all Calderón-Zygmund operators are bounded on $M^{p,q}_{w}(\mathbf{R}^{d};\mathcal{P}_{0})$ if and only if
@@
Observing that
$M(\operatorname{\mathbf{1}}_{Q})(x)\eqsim_{d}\frac{|Q|}{(\ell(Q)+|x-c_{Q}|)^{d}},$
where $\ell(Q)$ and $c_{Q}$ respectively denote the side length and the center of $Q$, one can recognize (5.2) as a condition appearing in two-weight boundedness characterizations of singular integrals from, e.g., the work of Muckenhoupt and Wheeden [MW76] and subsequent works.
Combining (5.2) with Theorem 1.2, we obtain the following characterization.
@@
Let $\mathcal{P}_{0}$ denote the collection of cubes centered at 0 and let $1<p\leq q<\infty$. Then the following are equivalent:
@@ Appendix A Weak-type bounds for sparse operators
Let $\mathcal{D}$ be a dyadic grid, let $0<\eta<1$, and let $\mathcal{S}\subseteq\mathcal{D}$ be a finite $\eta$-sparse collection. Then for all $0<\nu<1$ and all $f\in L^{1}_{\text{loc}}(\mathbf{R}^{d})$ there exists a collection $\mathcal{E}\subseteq\mathcal{D}$ such that
@@
[…]
@@
[…]
@@
Fix $Q\in\text{ch}_{\mathcal{E}}(Q_{0})$ and pick $\E\widehat{x}\in\widehat{Q}\backslash$ E. Then
@@
Let $\mathcal{D}$ be a dyadic grid, let $0<\nu<1$, and let $\mathcal{S}\subseteq\mathcal{D}$ be a finite collection satisfying
@@
[…]
@@
[…]
@@
[…]